Anna and Leah, living 39 km apart, leave their homes at the same time riding motorbikes and travel toward each other. Anna's average speed is 2 km per hour more than Leah's average speed. If they meet in 90 minutes, what is the average speed of Anna?
Answer by richwmiller(17219)

1.5x+1.5(x+2)=39
3/2x+3/2(x+2)=39
3x+3(x+2)=78
3x+3x+6=78
6x=72
x=12 Leah
x+2=14 Anna
check
1.5*12+1.5*14=39
18+21=39
39=39
ok